OpenStudy (anonymous):

How many solutions does this system have. -3x + 6y= 10 and -3x + 6y=-4 None, one, infinite or two? I think it's infinite but not sure

OpenStudy (campbell_st):

the lines are parallel.. so don't intersect... hence no solutions write them in slope intercept for and its y = 1/2 x + 10/6 and y = 1/2 x - 4/6 hope this helps

OpenStudy (anonymous):

Thanks, it does. I will rework to get my x and y.

OpenStudy (campbell_st):

you don't have to rework anything...other than showing there are parallel....same slope different y- intercepts..

OpenStudy (anonymous):

Also you can see that there is no solution from inspection, that is, the same expression is on the LHS of both equations but two different expressions are on the RHS. Also you can fint the determinant of the matrix \[\left[\begin{matrix}-3 & 6 \\ -3 & 6\end{matrix}\right]\] and see that it is equal to zero which means 1 of two things, either there is an infinitely many or no solution.